AM3 -Esercitazione 1 A.A. 2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 27 febbraio 2004

Esercizi

Calcolare i seguenti integrali indefiniti:

1. ∫ \frac{1 + 2e^{x}}{e^{2x} − 1} dx

2. ∫ \frac{1 + cos x}{1 − cos x} dx

3. ∫ x(1 + x^{2})e^{x^{2}} ln x dx

4. ∫ \frac{dx}{(x + 1)^{2}\sqrt{x^{2} + 2x + 2}}

5. ∫ \sqrt{x^{2} − 4} dx